From GlueXWiki
Jump to: navigation, search

May 29, 2014 FDC+CDC meeting


  1. Gas system gas panel
  2. DAQ (Beni)
  3. CDC status (Mike)
  4. Engineering (Bill)
  5. Electronics (Chris, Nick)
  6. FDC alignment [1] (Lubomir)
  7. Other

Instructions for Bluejeans meeting connection

FDC meeting ID: 290664653


Participants: Curtis, Naomi (CMU), Eugene, Bill, Dave, Chris, Nick, Simon, Fernando, Mike, and Lubomir (JLab).

Gas system

- Both CDC and FDC are now connected to the gas system. Some minor issues (like tube/cable swapping, gas settings) were resolved. FDC bubbles at 50 sccpm per package but now we are flushing it with 100 sccpm (nominal is 200 sccpm per package). CDC is being flushed with 1l/m (nominal 3l/m). From the gas GUI for experts (attached) we see the pressures in the individual cells within a FDC package are very similar. Bill is surprised and requested testing at different scanning times. When all the solenoid valves are closed the pressure sensor reads high pressure that increases; could be explained by heat flow.

- We were having problems with the gas control panel for the FDC: the display would disappear and then it requires turning off/on the power several times to reappear and some of the settings may be wrong. We are also not sure if it is only display problem or also the gas flow is disrupted. Maybe related to that: we need to fill the mixing tank with a flow that is a factor of 2 bigger than what is expected from the FDC usage. We need a spare control panel (Eugene requested a price estimate).

- It turned out the FDC part, including the mixing tank, were flushed over the weekend with 60/40 Ar/CO2 gas mixture, instead of 40/60. On Tuesday/Wednesday the detector was also flushed with the wrong gas mixture and as a result there were HV trips at that time.

CDC status

- Mike: still can't have all the HV channels working: the problems is in the CAEN modules. Will use one 1535P module instead of 1550P but that is not enough. Mike will do another CDC leakage test. The temperatures at the pre-amps are oscillating by 1-1.5 deg; still using the same cooling.


- Fernando: will discuss the HV issues with the CAEN people and request additional checks. Two HV modules are being repaired now, three were repaired before, all having problems with setting/holding HV. In addition we have many channels showing dark current of 0.5-1 uA without anything connected to the module. Eugene wanted to know how many spare modules we will need.


- We discussed once more if we want to scan the spare cathode with the JLab CMM machine; Eugene is reluctant since it requires (Bill) special stiff arm to hold the camera. Instead Eugene suggested to do measurements with a special ruler/scale (Fernando has it) only at the places where we suspect the original scan had problems.

- The chiller was put on scale and will check the weight in a month to estimate the Fluorinert leakage. Casey got the parts to make the extension for the sniffer to look for the leakage.

FDC alignment

- Lubomir showed results (linked above) about the FDC resolutions. Applying the offsets and rotations shown in the table, the cosmic tracks were reconstructed using cells #1,2, and 4, and the resolutions in x (from wires) and in y (from strips) for cell #3 were estimated. The wire resolution over the whole cell is ~150 microns in the 1-4mm region, which is a "record-high" resolution for the FDC and a result of the precise alignment. The strip resolution is 220 microns while we expect it to be ~150-200 microns, which means better strip alignment is needed.